1. A vendor prepares 100.00 hotdogs every day and sells at $19.00 /piece. For each hot dog, he spends $12.00 in the raw material. Additionally he spends $1.97 for packing each hotdog and monthly $46.00, $20.00, $9.00 as food truck rent, electricity and other expenses respectively. On a particular day in June 114.00 people came wanting the hotdog. Vendor thinks $3.00 per lost sale. Determine vendor’s profit for that day? Assume there are 30 days in the month.

2.For a table manufacturing company, selling price for a table is $65.00 per Unit, Variable cost is $15.00 per Unit, labor charge is $10.44 per Unit, rent is $573.00 per month and transportation is $10 per 10 tables. If 50 tables are sold in a month how much is the fixed cost for that month?

please, show processes to solve them
Answer #1

Answer No 1

Particular Amount($)
SP /Piece 19
Raw Material 12
Packing 1.97
Contribution/Piece 5.03
Units prepared /day 100
Total Contribution 503
Loss on Lost sale(14 pieces*3) 42
Total Income/Day 461

Notes-

1-FIxed cost being sunk cost has not been considered while making a decision.

2- Loss on lost sale has been considered while calculating income because Its loss due to less production of hot dogs.

Answer -2

Particular Amount($)
SP /Piece 65
VC/Unit 15
Labour Charges/unit 10.44
Contribution/Piece 39.56
Units Sold During the month 50
Total Contribution 1978

Rent/month=$573

Transportation cost is $10 for 10 tables so 50 tables are sold so transportation cost for the month=10*5=$50

Total fixed cost=rent+transportation cost=573+50=$623.